# Facility Location Games with Scaling Effects

### v1

@@ 6 Approximation Ratio Results
$\max\limits_{q\in\mathcal{Q},\boldsymbol{x}\in X^{n}}\frac{TC(q,f(\boldsymbol{x}),\boldsymbol{x})}{TC(q,y^{*}_{TC},\boldsymbol{x})}.$
@@
For simplicity of notation, we let $r_{q}:=\frac{\max_{y\in X}q(y)}{\min_{y\in X}q(y)}$ be the ratio between the highest and lowest values of the scaling function.
We begin with an impossibility result in the general continuous setting, which explains our need to restrict the scaling function parameters to ensure single-peakedness and widen the space of strategyproof and anonymous mechanisms.
Theorem 6.1
Every strategyproof and anonymous mechanism has an unbounded approximation ratio for total and maximum cost.
Proof
By Moulin [17]’s characterization, a strategyproof and anonymous mechanism in our setting must be a phantom mechanism, as the classic setting is a special case of our setting. By Theorem 5.1, we also know that a phantom mechanism is only strategyproof if all $n+1$ phantoms are located at the same location. Such a phantom mechanism is equivalent to the constant mechanism, which has an unbounded approximation ratio as it is not unanimous.
As we will show, our approximation ratio results are typically a function of $r_{q}$. We show that when the scaling function is chosen such that agents have single-peaked preferences, the value of $r_{q}$ is upper bounded by a constant.
@@
Lastly, consider the intervals where $q^{\prime}(y)<0$, and consequently $q$ satisfies $q(y)+q^{\prime}(y)\geq 0$. Similarly, we define $g(y):=e^{-y}q(y)$, where $g^{\prime}(y)=e^{-y}(q^{\prime}(y)-q(y))$. We know $g(y)$ is non-increasing because $q^{\prime}(y)-q(y)\leq 0$. Consider the points $y_{1},y_{2}\in[0,1]$ where $y_{1}<y_{2}$. Since $g(y_{1})\geq g(y_{2})$, we know that $e^{-y_{1}}q(y_{1})\geq e^{-y_{2}}q(y_{2})$. Also, since $q(y)>0$ and $e^{y}>0$, we have $\frac{q(y_{2})}{q(y_{1})}\leq e^{y_{2}-y_{1}}$, which is again maximized when $y_{1}=0$ and $y_{2}=1$. Therefore $r_{q}\leq e$.
@@
Suppose for $j\in\{1,\dots,k\}$ that each line segment of $q$ is denoted as $q_{j}(y)=a_{j}y+b_{j}$. From Theorem 5.3, for agents to always have single-peaked preferences, we require that $q(y)-|q^{\prime}(y)|\geq 0$ for all $y\in[0,1]\backslash\{D\}$, where $\{D\}$ denotes the set of points where $q(y)$ is not differentiable. This implies that for each line segment, if $a_{j}>0$, then $a_{j}y+b_{j}-a_{j}>0\implies b_{j}\geq(1-y)a_{j}$. Similarly, if $a_{j}<0$, then $b_{j}\geq-(1+y)a_{j}$.
@@
If $a_{j}>0$, then $r_{q_{j}}=\frac{q_{j}(s_{j})}{q_{j}(s_{j-1})}=\frac{a_{j}s_{j}+b_{j}}{a_{j}s_{j-1}+b_{j}}$. Since we require $b_{j}\geq(1-y)a_{j}$ on $y\in[s_{j-1},s_{j}]$, we know that $b_{j}\geq(1-s_{j-1})a_{j}$. Therefore $r_{q_{j}}=\frac{a_{j}s_{j}+b_{j}}{a_{j}s_{j-1}+b_{j}}\leq\frac{a_{j}s_{j}+(1-s_{j-1})a_{j}}{a_{j}s_{j-1}+(1-s_{j-1})a_{j}}=(1+s_{j}-s_{j-1})$.
If $a_{j}<0$, then $r_{q_{j}}=\frac{q_{j}(s_{j-1})}{q_{j}(s_{j})}=\frac{a_{j}s_{j-1}+b_{j}}{a_{j}s_{j}+b_{j}}$. Since we require $b_{j}\geq-(1+y)a_{j}$ on $y\in[s_{j-1},s_{j}]$, we know that $b_{j}\geq-(1+s_{j})a_{j}$. Therefore $r_{q_{j}}=\frac{a_{j}s_{j-1}+b_{j}}{a_{j}s_{j}+b_{j}}\leq\frac{a_{j}s_{j-1}-(1+s_{j})a_{j}}{a_{j}s_{j}-(1+s_{j})a_{j}}=(1+s_{j}-s_{j-1})$, and we know that for each $q_{j}$, $r_{q_{j}}\leq(1+s_{j}-s_{j-1})$.
Since $q$ is piecewise linear, its global maxima and minima must be elements of $S$. We take an arbitrary global minimum $s_{min}:=\arg\min_{y\in X]}q(y)$ and global maximum $s_{max}:=\arg\max_{y\in X}q(y)$, where $s_{min},s_{max}\in S$, and suppose w.l.o.g. that $s_{min}<s_{max}$. It also suffices to assume that for each line segment $q_{j}$ between $s_{min}$ and $s_{max}$, we have $a_{j}>0$ (i.e. the function is strictly increasing in $[s_{min},s_{max}]$). This means that $r_{q}$ is the product of the $r_{q_{j}}$ between $s_{min}$ and $s_{max}$. Finally, since $1\leq r_{q_{j}}\leq(1+s_{j}-s_{j-1})$, and applying the arithmetic mean-geometric mean inequality, we have $r_{q}\leq\prod\limits_{i=1}^{k}(1+s_{i}-s_{i-1})\leq(1+\frac{1}{k})^{k}$.
@@
We follow up with another result which restricts the space of phantom mechanisms with a bounded approximation ratio.
@@
A phantom mechanism which does not place at least one phantom on 00 and $1$ has an unbounded approximation ratio for total cost and maximum cost.
@@
Due to symmetry it suffices to consider phantom mechanisms which do not have any phantoms on $1$. If all agents are located at $1$, the optimal total and maximum cost are 00, but such a phantom mechanism must place the facility at a location strictly left of $1$, as all $n+1$ phantoms are strictly left of $1$. Therefore the total cost and maximum cost approximation ratios are unbounded.
This lemma shows that when proving the approximation results, it suffices to consider phantom mechanisms where $p_{1}=0$ and $p_{n+1}=1$. Such a phantom mechanism is equivalent to a phantom mechanism with only $n-1$ phantoms, and is known to characterize the strategyproof and anonymous mechanisms which place the facility between the leftmost and rightmost agent locations [17].
@@ 6.1 Total Cost
We now prove approximation ratio results for the objective of total cost, a standard utilitarian measure of efficiency. We begin by giving a lower bound on the approximation ratio achievable by a phantom mechanism, with a proof which applies to both continuous and piecewise linear scaling functions.
@@
For continuous and piecewise linear scaling functions, any phantom mechanism has a total cost approximation ratio of at least $r_{q}$.
@@
We consider two cases: either all phantoms are located at 00 or $1$, or there is at least one phantom that is not located at 00 or $1$. We start with the former case, where the facility is located at some agent. For every such phantom mechanism, we can always construct an instance where the scaling function is highest at the agent that receives the facility, so the corresponding total cost must be at least $\max_{y}q(y)\sum_{i}|f_{med}(\boldsymbol{x})-x_{i}|$, where $f_{med}$ is the median mechanism. Now suppose that the instance has an even number of agents, with $x_{n/2}\neq x_{n/2+1}$, and that the scaling function is lowest at $(x_{n/2}+x_{n/2+1})/2$. Placing the facility at this point results in the optimal total cost of $\min_{y}q(y)\sum_{i}|f_{med}(\boldsymbol{x})-x_{i}|$, and thus the total cost approximation ratio must be at least $r_{q}$ in this case.
Now consider the case where at least one phantom is located in $(0,1)$. By Lemma 1, it suffices to consider phantom mechanisms with at least one phantom at 00 and $1$, so we are interested in where the remaining $n-1$ phantoms are placed. We show that for every phantom mechanism $f$ where at least one phantom $p_{j}$ on $(0,1)$ is not at the lowest point of the scaling function, there exists a phantom mechanism $f^{\prime}$ that places the $n-1$ phantoms at the lowest point of the scaling function and has a strictly lower approximation ratio. Specifically, if the scaling function has multiple global minima, the phantoms are placed at the leftmost global minimum on $(0,1)$. If the scaling function only has global minima on 00 and/or $1$, then for sufficiently small $\epsilon>0$, $f^{\prime}$ places the phantoms at $\epsilon$ or $1-\epsilon$. For every agent location profile $\boldsymbol{x}$ such that $f(\boldsymbol{x})=p_{j}$, we claim there exists another profile $\boldsymbol{x}^{\prime}$ where, under $f^{\prime}$, the total cost ratio with the optimal total cost is strictly lower. We derive $\boldsymbol{x}^{\prime}$ from $\boldsymbol{x}$ by first moving each agent closer to $f(\boldsymbol{x})=p_{j}$ by a sufficiently small constant factor $c>0$ (so that each agent $i$’s distance changes from $|x_{i}-p_{j}|$ to $c|x_{i}-p_{j}|$), and then translating the profile such that the facility is co-located [^4] This occurs as long as there are agents on both sides of the phantoms. Since the phantoms are placed at $\epsilon$ and $1-\epsilon$ instead of 00 and $1$, there is space on both sides of the phantoms. with the $n-1$ phantoms of $f^{\prime}$ and the relative distances between the agents and the facility is the same but multiplied by the sufficiently small constant factor. The distances are still $c|x_{i}-p_{j}|$, so the ratio of total unscaled distances is still the same. But now the facility is placed at a point where the scaling function value is strictly lower, so the total cost ratio is lower. Therefore $f^{\prime}$ has the lowest approximation ratio of all phantom mechanisms which have at least one phantom on $(0,1)$.
@@
We now show that $f^{\prime}$ has an approximation ratio of at least $r_{q}$. We define a scaling function $q$ which has a unique global minimum at some point $c_{1}\in(0,1)$, a unique global maximum at $c_{2}>c_{1}$, and a point $c_{3}>c_{2}$ where $q(c_{3})=q(c_{1})+\epsilon$, for sufficiently small $\epsilon>0$. Consider a location profile where $x_{1}=c_{2}<x_{2}=x_{3}=\dots=x_{n}=c_{3}$. The mechanism $f^{\prime}$ places the facility at $x_{1}$, resulting in a total cost of $(n-1)(x_{n}-x_{1})\max_{y}q(y)$, whilst the optimal total cost is at most $(x_{n}-x_{1})(\min_{y}q(y)+\epsilon)$ by placing the facility at $c_{3}$. Therefore $f^{\prime}$ has an approximation ratio of at least $r_{q}$. By exhaustion of cases, we see that any phantom mechanism has a total cost approximation ratio of at least $r_{q}$.
@@
For continuous and piecewise linear scaling functions, the median mechanism has a total cost approximation ratio of $r_{q}$.
@@
The median mechanism minimizes the sum of distances, so its total cost is at most $\max_{y}q(y)\sum_{i}|f_{med}(x)-x_{i}|$, whilst the optimal total cost is at least $\min_{y}q(y)\sum_{i}|f_{med}(x)-x_{i}|$. This shows the approximation ratio is at most $r_{q}$, and we have a matching lower bound from Theorem 6.4.
Recall that phantom mechanisms characterize all strategyproof and anonymous mechanisms when agents have single-peaked preferences. We can consider scaling functions which ensure that agents have single-peaked preferences, and use Theorems 6.2 and 6.3 to find the specific total cost approximation ratio lower bounds for any strategyproof and anonymous mechanism.
@@
For continuous scaling functions, when agents are guaranteed to have single-peaked preferences, any strategyproof and anonymous mechanism has a total cost approximation ratio of at least $e$.
@@
Again, the median mechanism would also give the best total cost approximation ratio, and can be suitably implemented for these restricted sets of scaling functions. However, in a scenario where the scaling function does not guarantee single-peaked agent preferences, phantom mechanisms no longer guarantee strategyproofness, and a constant mechanism has an unbounded approximation ratio. The dictator mechanism, which selects a specific agent to be the dictator and places the facility at their reported location, is known to be strategyproof for general agent preferences (though it is clearly not anonymous). We show that this mechanism has a bounded approximation ratio for total cost.
@@
Suppose there are $n-1$ agents at the same location away from the dictator. Also suppose that the scaling function is highest at the dictator’s location, and lowest at the other $n-1$ agents’ location. The total cost approximation ratio is therefore at least $(n-1)r_{q}$. To see that the approximation ratio cannot be higher than this, note that the total cost corresponding to the dictator mechanism is at most $d(n-1)\max_{y\in X}q(y)$ where $d\in(0,1]$, as the dictator receives 00 cost, and the optimal total cost is at least $d\min_{y\in X}q(y)$ when not all agents are at the same point.

### v2

@@ 7 Approximation Ratio Results
$\max\limits_{q\in\mathcal{Q},\boldsymbol{x}\in X^{n}}\frac{TC(q,f(q,\boldsymbol{x}),\boldsymbol{x})}{TC(q,y^{*}_{TC},\boldsymbol{x})}.$
@@
For simplicity of notation, we let $r_{q}:=\frac{\max_{y\in X}q(y)}{\min_{y\in X}q(y)}$ be the ratio between the highest and lowest values of the scaling function. As we will show, our approximation ratio results are typically a function of $r_{q}$. We show that when the scaling function is chosen such that agents have single-peaked preferences, the value of $r_{q}$ is upper bounded by a constant.
@@
Lastly, consider the intervals where $q^{\prime}(y)<0$, and consequently $q$ satisfies $q(y)+q^{\prime}(y)\geq 0$. Similarly, we define $g(y):=e^{-y}q(y)$, where $g^{\prime}(y)=e^{-y}(q^{\prime}(y)-q(y))$. We know $g(y)$ is non-increasing because $q^{\prime}(y)-q(y)\leq 0$. Consider the points $y_{1},y_{2}\in[0,1]$ where $y_{1}<y_{2}$. Since $g(y_{1})\geq g(y_{2})$, we know that $e^{-y_{1}}q(y_{1})\geq e^{-y_{2}}q(y_{2})$. Also, since $q(y)>0$ and $e^{y}>0$, we have $\frac{q(y_{2})}{q(y_{1})}\leq e^{y_{2}-y_{1}}$, which is again maximized when $y_{1}=0$ and $y_{2}=1$. Therefore $r_{q}\leq e$. ∎
@@
Suppose for $j\in\{1,\dots,k\}$ that each line segment of $q$ is denoted as $q_{j}(y)=a_{j}y+b_{j}$.
From Theorem 5.3, for agents to always have single-peaked preferences, we require that $q(y)-|q^{\prime}(y)|\geq 0$ for all $y\in[0,1]\backslash\{D\}$, where $\{D\}$ denotes the set of points where $q(y)$ is not differentiable. This implies that for each line segment, if $a_{j}>0$, then $a_{j}y+b_{j}-a_{j}>0\implies b_{j}\geq(1-y)a_{j}$. Similarly, if $a_{j}<0$, then $b_{j}\geq-(1+y)a_{j}$.
@@
If $a_{j}>0$, then
$r_{q_{j}}=\frac{q_{j}(s_{j})}{q_{j}(s_{j-1})}=\frac{a_{j}s_{j}+b_{j}}{a_{j}s_{j-1}+b_{j}}.$
Since we require $b_{j}\geq(1-y)a_{j}$ on $y\in[s_{j-1},s_{j}]$, we know that $b_{j}\geq(1-s_{j-1})a_{j}$. Therefore
$r_{q_{j}}=\frac{a_{j}s_{j}+b_{j}}{a_{j}s_{j-1}+b_{j}}\leq\frac{a_{j}s_{j}+(1-s_{j-1})a_{j}}{a_{j}s_{j-1}+(1-s_{j-1})a_{j}}=(1+s_{j}-s_{j-1}).$
If $a_{j}<0$, then
$r_{q_{j}}=\frac{q_{j}(s_{j-1})}{q_{j}(s_{j})}=\frac{a_{j}s_{j-1}+b_{j}}{a_{j}s_{j}+b_{j}}.$
Since we require $b_{j}\geq-(1+y)a_{j}$ on $y\in[s_{j-1},s_{j}]$, we know that $b_{j}\geq-(1+s_{j})a_{j}$. Therefore
$r_{q_{j}}=\frac{a_{j}s_{j-1}+b_{j}}{a_{j}s_{j}+b_{j}}\leq\frac{a_{j}s_{j-1}-(1+s_{j})a_{j}}{a_{j}s_{j}-(1+s_{j})a_{j}}=(1+s_{j}-s_{j-1}),$
and we know that for each $q_{j}$, $r_{q_{j}}\leq(1+s_{j}-s_{j-1})$.
Since $q$ is piecewise linear, its global maxima and minima must be elements of $S$. We take an arbitrary global minimum $s_{min}:=\arg\min_{y\in X]}q(y)$ and global maximum $s_{max}:=\arg\max_{y\in X}q(y)$, where $s_{min},s_{max}\in S$, and suppose w.l.o.g. that $s_{min}<s_{max}$. It also suffices to assume that for each line segment $q_{j}$ between $s_{min}$ and $s_{max}$, we have $a_{j}>0$ (i.e. the function is strictly increasing in $[s_{min},s_{max}]$). This means that $r_{q}$ is the product of the $r_{q_{j}}$ between $s_{min}$ and $s_{max}$.
Finally, since $1\leq r_{q_{j}}\leq(1+s_{j}-s_{j-1})$, and applying the arithmetic mean-geometric mean inequality, we have $r_{q}\leq\prod\limits_{i=1}^{k}(1+s_{i}-s_{i-1})\leq(1+\frac{1}{k})^{k}$. ∎
@@
We follow up with another result which restricts the space of phantom mechanisms with scaling when we require a bounded approximation ratio.
@@
Let $f$ be a phantom mechanism with scaling, with phantoms $p_{1}(q),\dots,p_{n+1}(q)$. If there exists a scaling function $q$ such that $p_{i}(q)\neq 0$ and/or $p_{i}(q)\neq 1$ for all $i\in[n+1]$, then $f$ has an unbounded approximation ratio for total cost and maximum cost.
@@
Due to symmetry it suffices to consider a scaling function $q$ such that $p_{i}(q)\neq 1$ for all $i\in[n+1]$. If all agents are located at $1$, the optimal total and maximum cost are 00, but $f$ places the facility at a location strictly left of $1$, as all $n+1$ phantoms are strictly left of $1$. Therefore the total cost and maximum cost approximation ratios are unbounded. ∎
This lemma shows that when proving the approximation results, it suffices to consider phantom mechanisms with scaling where $p_{1}=0$ and $p_{n+1}=1$. This is equivalent to a phantom mechanism with scaling where there are only $n-1$ phantoms, and by Theorem 6.1, this characterizes the strategyproof and anonymous mechanisms which place the facility between the leftmost and rightmost agent locations.
We next prove a lemma which, given a scaling function $q$, characterizes phantom mechanisms with scaling that have the best approximation ratio among all such mechanisms which have at least one phantom in $(0,1)$. This statement holds for both objectives of total and maximum cost.
Lemma 2
Let $f$ be a phantom mechanism with scaling, where $p_{1}=0$ and $p_{n+1}=1$. For any given continuous scaling function $q$ with a global minimum in $(0,1)$, if $p_{i}\in(0,1)$ for some $i\in\{2,\dots,n\}$, then there exists a phantom mechanism with scaling $f^{\prime}$ that places the $n-1$ phantoms at a global minimum of the scaling function and has a weakly lower approximation ratio.
Furthermore, for any given continuous scaling function $q$ which only has global minima in $\{0,1\}$, if $p_{i}\in(0,1)$ for some $i\in\{2,\dots,n\}$, then for sufficiently small $\epsilon>0$, there exists a phantom mechanism with scaling $f^{\prime}$ that places the $n-1$ phantoms at $\epsilon$ or $1-\epsilon$ and has a weakly lower approximation ratio.
Proof
Throughout the proof, we are primarily concerned with the placement of the $n-1$ phantoms $p_{2},\dots,p_{n}$. We show that for every phantom mechanism with scaling $f$ where at least one phantom $p_{j}$ on $(0,1)$ is not at the lowest point of the scaling function, there exists a different phantom mechanism with scaling $f^{\prime}$ that has a strictly lower approximation ratio. This mechanism places the $n-1$ phantoms at the global minimum of $q$ if it lies in $(0,1)$. Specifically, if the scaling function has multiple global minima, the phantoms are placed at the leftmost global minimum on $(0,1)$. If the scaling function only has global minima on 00 and/or $1$, then for sufficiently small $\epsilon>0$, $f^{\prime}$ places the phantoms at $\epsilon$ or $1-\epsilon$.
For every agent location profile $\boldsymbol{x}$ such that $f(\boldsymbol{x})=p_{j}$, we claim there exists another profile $\boldsymbol{x}^{\prime}$ where, under $f^{\prime}$, the total cost ratio with the optimal total cost is strictly lower. We derive $\boldsymbol{x}^{\prime}$ from $\boldsymbol{x}$ by first moving each agent closer to $f(\boldsymbol{x})=p_{j}$ by a sufficiently small constant factor $c>0$ (so that each agent $i$’s distance changes from $|x_{i}-p_{j}|$ to $c|x_{i}-p_{j}|$), and then translating the profile such that the facility is co-located [^5] This occurs as long as there are agents on both sides of the phantoms. Since the phantoms are placed at $\epsilon$ and $1-\epsilon$ instead of 00 and $1$, there is space on both sides of the phantoms. with the $n-1$ phantoms of $f^{\prime}$ and the relative distances between the agents and the facility is the same but multiplied by the sufficiently small constant factor. The distances are still $c|x_{i}-p_{j}|$, so the ratio of total/maximum unscaled distances is still the same. But now the facility is placed at a point where the scaling function value is strictly lower, so the total/maximum cost ratio is lower. Therefore $f^{\prime}$ has the lowest approximation ratio of all phantom mechanisms with scaling which have at least one phantom on $(0,1)$. ∎
@@ 7.1 Total Cost
We now prove approximation ratio results for the objective of total cost, a standard utilitarian measure of efficiency. We begin by giving a lower bound on the approximation ratio achievable by a phantom mechanism, with a proof which applies to any continuous scaling function.
@@
Any phantom mechanism with scaling has a total cost approximation ratio of at least $r_{q}$.
@@
Given a scaling function $q$, we consider two cases: either all phantoms are located at 00 or $1$, or there is at least one phantom that is not located at 00 or $1$. We start with the former case, where the facility is located at some agent. For every such phantom mechanism with scaling, we can always construct an agent location profile where the scaling function is highest at the agent that receives the facility, so the corresponding total cost must be at least $\max_{y}q(y)\sum_{i}|f_{med}(\boldsymbol{x})-x_{i}|$, where $f_{med}$ is the median mechanism. Now suppose that the agent location profile has an even number of agents, with $x_{n/2}\neq x_{n/2+1}$, and that the location profile is constructed such that the scaling function is lowest at $(x_{n/2}+x_{n/2+1})/2$. Placing the facility at this point results in the optimal total cost of $\min_{y}q(y)\sum_{i}|f_{med}(\boldsymbol{x})-x_{i}|$, and thus the total cost approximation ratio must be at least $r_{q}$ in this case.
Now consider the case where at least one phantom is located in $(0,1)$. From Lemma 2, it suffices to consider mechanism $f^{\prime}$, a phantom mechanism with scaling which places one phantom on 00, one phantom on $1$, and the remaining $n-1$ phantoms on the leftmost global minimum of $q$ if there exists a global minimum of $q$ in $(0,1)$. If all global minima of $q$ are in $\{0,1\}$, then for sufficiently small $\epsilon>0$, the $n-1$ phantoms are placed on $\epsilon$ or $1-\epsilon$.
@@
We show that $f^{\prime}$ has an approximation ratio of at least $r_{q}$. We define a scaling function $q$ which has a unique global minimum at some point $c_{1}\in(0,1)$, a unique global maximum at $c_{2}>c_{1}$, and a point $c_{3}>c_{2}$ where $q(c_{3})=q(c_{1})+\epsilon$, for sufficiently small $\epsilon>0$. Consider a location profile where $x_{1}=c_{2}<x_{2}=x_{3}=\dots=x_{n}=c_{3}$. The mechanism $f^{\prime}$ places the facility at $x_{1}$, resulting in a total cost of $(n-1)(x_{n}-x_{1})\max_{y}q(y)$, whilst the optimal total cost is at most $(x_{n}-x_{1})(\min_{y}q(y)+\epsilon)$ by placing the facility at $c_{3}$. Therefore $f^{\prime}$ has an approximation ratio of at least $r_{q}$. By exhaustion of cases, we see that any phantom mechanism has a total cost approximation ratio of at least $r_{q}$. ∎
@@
The median mechanism has a total cost approximation ratio of $r_{q}$.
@@
The median mechanism minimizes the sum of (unscaled) distances, so its total cost is at most $\max_{y}q(y)\sum_{i}|f_{med}(x)-x_{i}|$, whilst the optimal total cost is at least $\min_{y}q(y)\sum_{i}|f_{med}(x)-x_{i}|$. This shows the approximation ratio is at most $r_{q}$, and we have a matching lower bound from Theorem 7.3. ∎
Recall that phantom mechanisms with scaling characterize all strategyproof and anonymous mechanisms when agents have single-peaked preferences. We can consider scaling functions which ensure that agents have single-peaked preferences, and use Theorems 7.1 and 7.2 to find the specific total cost approximation ratio lower bounds for any strategyproof and anonymous mechanism.
@@
For arbitrary continuous scaling functions, when agents are guaranteed to have single-peaked preferences, any strategyproof and anonymous mechanism has a total cost approximation ratio of at least $e$.
@@
Again, the median mechanism would also give the best total cost approximation ratio, and can be suitably implemented for these restricted sets of scaling functions. However, in a scenario where the scaling function does not guarantee single-peaked agent preferences, phantom mechanisms with scaling no longer guarantee strategyproofness, and a constant mechanism has an unbounded approximation ratio. The dictator mechanism, which selects a specific agent to be the dictator and places the facility at their reported location, is known to be strategyproof for general agent preferences (though it is clearly not anonymous). We show that this mechanism has a bounded approximation ratio for total cost.
@@
Suppose there are $n-1$ agents at the same location away from the dictator. Also suppose that the scaling function is highest at the dictator’s location, and lowest at the other $n-1$ agents’ location. The total cost approximation ratio is therefore at least $(n-1)r_{q}$. To see that the approximation ratio cannot be higher than this, note that the total cost corresponding to the dictator mechanism is at most $d(n-1)\max_{y\in X}q(y)$ where $d\in(0,1]$, as the dictator receives 00 cost, and the optimal total cost is at least $d\min_{y\in X}q(y)$ when not all agents are at the same point. ∎